Samuraiduck:

For the equation y = 3x2 − 4x + 11, choose the correct application of the quadratic formula.

Echofire:

What are the options?

Samuraiduck:

oh sorry they are A= x equals four plus or minus the square root of negative four squared minus four times three times eleven, all divided by two times three b=x equals negative four plus or minus the square root of negative four squared minus four times three times eleven, all divided by two times three C=x equals negative three plus or minus the square root of three squared minus four times negative four times eleven, all divided by two times negative four D=x equals three plus or minus the square root of three squared minus four times negative four times eleven, all divided by two times negative four.

Samuraiduck:

sorry the answers are photos so its hard to copy

Renne:

Send screen shots of the photos please

Echofire:

I suggest using Symbolab to type it up and save images of it. You can also type it on Symbolab and post a link here.

Echofire:

1 attachment
Renne:

So 11+4= 15- 3 to the 2 power which would be 6

Samuraiduck:

1 attachment
Samuraiduck:

there it is

Renne:

Okay lets see-

Echofire:

First, utilize this:

Samuraiduck:

ohh thank you

Echofire:

@samuraiduck wrote:
ohh thank you
You are welcome. : )

Renne:

To find the roots of the equation y = 3x2 − 4x + 11, we can use the quadratic formula. The formula is (-b ± √(b² - 4ac)) / 2a, where a, b, and c are the coefficients of the quadratic equation ax² + bx + c = 0. In this case, a = 3, b = -4, and c = 11. Substituting these values into the formula, we get (-(-4) ± √((-4)² - 4(3)(11))) / 2(3). Simplifying this expression, we get (4 ± √(16 - 132)) / 6. Further simplifying, we get (4 ± √(-116)) / 6. Since the square root of a negative number is not a real number, the equation has no real roots.

XLynx:

@renne wrote:
To find the roots of the equation y = 3x2 − 4x + 11, we can use the quadratic formula. The formula is (-b ± √(b² - 4ac)) / 2a, where a, b, and c are the coefficients of the quadratic equation ax² + bx + c = 0. In this case, a = 3, b = -4, and c = 11. Substituting these values into the formula, we get (-(-4) ± √((-4)² - 4(3)(11))) / 2(3). Simplifying this expression, we get (4 ± √(16 - 132)) / 6. Further simplifying, we get (4 ± √(-116)) / 6. Since the square root of a negative number is not a real number, the equation has no real roots.
You could just calculate if the discriminant is negative or positive. Seems like you just took 5 more steps than I did to arrive at the same answer.